# shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation

### v1

@@ 4 Taxonomy
We note that, depending on the explainability method, these scores might be inverted, following the interpretation “errors are not important for a metric to achieve a high score”.121212Most gradient and attention based approaches return high scores for errors (e.g. Treviso et al. 2021; Rubino et al. 2021), for embedding-based approaches it depends on whether cosine-similarity or -distance is used (e.g. Leiter 2021; Tao et al. 2022), for attribution based techniques using LIME (Ribeiro et al. 2016) or SHAP (Lundberg and Lee 2017), low scores are returned for errors (e.g., as part of the baseline implementations in Eval4NLP21: https://github.com/eval4nlp/SharedTask2021/tree/main/baselines). Two shared tasks for the explainabiltiy of reference-free metrics have originated from the setting in Fomicheva et al. (2022) so far (Fomicheva et al. 2021; Zerva et al. 2022). As the respective papers already give an overview of the participating methods, we will only briefly describe the methods and set our focus to describe the model access they require. We describe the model access levels of Figure 4 from left to right. Approaches that work directly on black box models are explored in Fomicheva et al. (2022); Eksi et al. (2021); Treviso et al. (2021). They leverage model agnostic explainers such as LIME (Ribeiro et al. 2016) and SHAP (Lundberg and Lee 2017). These methods check how a metric’s sentence level score changes when the input sentence is perturbed at selected positions. While the results correlate with the human annotations to some degree, techniques that use model access performed better in the shared tasks. Gradient based methods are explored by Fomicheva et al. (2022); Eksi et al. (2021); Treviso et al. (2021); Rei et al. (2022a). These papers use methods like Integrated Gradients (Sundararajan et al. 2017) to determine the effect each input token has on the output. Attention values perform especially good for Treviso et al. (2021); Rei et al. (2022b), leading to winning techniques for most language pairs in both shared tasks, i.e., producing the most plausible explanations. In specific, Treviso et al. (2021) succeed by scaling attention weights by the norm of value vectors and filtering for attention heads that have the highest scores in detecting errors. Rei et al. (2022b) build on their approach and (1) instead scale attention weights with the l2-norm of the gradient of value vectors and (2) designing an ensemble method for attention heads. In their final submission, they ensemble the best performing attention heads over all layers. Word-level contextualized embeddings are used by Leiter (2021); Tao et al. (2022); Azadi et al. (2022). Specifically, they leverage the word-level embedding distances that are inherently computed in token-matching metrics (see §2.1) and use them as importance scores. For model-specific works, Rubino et al. (2021) jointly fine-tune their metrics on the word-level and sentence level using synthetic data, to directly output the computed word-level scores.131313Rubino et al. (2021) won the constrained track of Eval4NLP21, which did not allow for word-level finetuning. Kabir and Carpuat (2021) use LIME with Transquest (Ranasinghe et al. 2020a) and merge the word-level importance scores with additional word-level scores obtained from a pure word-level model. After the shared tasks, Fernandes et al. (2022b) enhance explainability techniques (in case of MT they use a gradient technique) by adding learnable parameters and employing a student teacher mechanism to improve the correlation to human word-level scores. Rei et al. (2023) extend the work of Rei et al. (2022b) by using reference-based metrics (e.g. COMET) and importance scores of the references. They construct a data set from the WMT21 MQM annotations Freitag et al. (2021b) by assuming that feature attribution scores link to annotated error spans. Evaluations on this data set show that the winning explainability techniques of the shared task perform best and that references can improve the plausibility. Notably, they also show that explanations can help to detect critical errors in translations. Based on this, they employ the explanations for metric diagnosis, to test whether their metrics are more prone to ignoring certain types of critical errors. In their limitations, they note that the explainabiltiy techniques they use have shown to be faithful in some scenarios, but also note that they did not specifically test whether they are faithful for their architecture. Finally, we note that many more metrics can provide feature-importance like explanations by design. Similar to the base metrics for Leiter (2021); Tao et al. (2022); Azadi et al. (2022), most unsupervised matching metrics could be used to determine importance scores based on word- or embedding-similarities. Further, the per-word generation probabilities in text generation metrics like PRISM or BARTScore can also be interpreted as feature importance (they can also be interpreted as the confidence of the generation models; Fomicheva et al. 2020).
@@
In the example presented in Figure 5, high feature importance scores are assigned to words that were translated correctly, i.e. “Things/complicated” and lower scores indicate words that are wrong.
@@
Goals: The discussed feature importance approaches are tested for plausibility, so strong approaches could be used for metric accessibility/improvement and semi-automatic labeling. The usage for metric diagnosis is currently limited, as only samples are explained, and approaches have not been tested for faithfulness. The current approaches have not been used to detect social biases.
Audiences: The relation of low/high scores and translation errors is intuitive, such that these techniques could be used by non-expert and expert users. Metric developers could use the methods to create new data sets or to directly incorporate them into metrics Leiter et al. (2022b).
Limitations: For all approaches, it is not specifically tested whether the explanations are faithful.141414In different possible aspects (Jacovi and Goldberg 2020) Notably, Rei et al. (2023) consider that faithfulness is probably given in their case but note the missing tests as limitation.
Second, certain translation errors cannot be easily captured by highlighting specific words. For example, the word-level scores cannot represent cases where the MT fails to explicitly express a grammatical feature that is omitted in the source language but is required to be explicit in the target language (for example, the use of articles when translating from Russian into German).
Third, the explanation type does not provide correspondence between highlighted error words in the source and target language.
@@
Fine-grained errors (Row 3 in Figure 4) Some recent works detect errors on a word-level and leverage their number and type to compute a sentence-level score (Lu et al. 2022, 2023; Bao et al. 2023; Xu et al. 2023). This follows the ideas of the MQM annotation scheme (Lommel et al. 2014; Freitag et al. 2021a) in which human experts evaluate translations in a similar manner. Freitag et al. (2021a) have shown that this annotation scheme improves over previous annotation methods such as direct assessments (Graham et al. 2015). While the main goal of these works is to provide sentence-level scores, they are also explainable-by-design and the generated error-labels carry explanatory value (model-specific in Figure 4). In specific, BARTScore++ (Lu et al. 2022) distinguishes explicit errors that are easy to detect and implicit errors. They detect the explicit errors based on BART generation probabilities. Then they generate a sentence-level score based on an additional improved hypothesis sentence in multiple BARTScore calculations. Lu et al. (2023) also combine fine-grained errors into a score. They first request ChatGPT to generate explanations based on MQM annotations (erroneous words, error type and error majority). In a second step, they ask a model to calculate a score from these. They note the instabilities of ChatGPT, but find that their approach is strong on a system-level and generates sentence-level scores that are better than BERTScore. Xu et al. (2023) fine-tune InstructScore, a LLaMA model (Touvron et al. 2023) that generates a text containing error majority, error spans, error types (MQM) and an explanation for every error in an input sentence. The generated texts are similar to Lu et al. (2023) but add more details on each error. Similarly, they construct a score from the count and severity of errors. They obtain synthetic training data from GPT4 (OpenAI 2023a), analyze failure modes of InstructScore during training and later optimize the metric on real machine translation outputs on data sets with annotated ground truth scores. Their model shows state-of-the-art results for reference-free sentence-level evaluations. Further, the detailed error annotations can be used as fine-grained explanations. Differently to the previous approaches, Bao et al. (2023) fine-tune a seq2seq model to predict addition and omission errors on a word-level. In the example presented in Figure 5, a tag “sense” flags the wrong word sense of “became”. Notably, all fine-grained error approaches we describe are very recent, which is potentially caused by the adaptation of MQM in the WMT metrics shared tasks and the recent successes of LLMs.
Goals: These fine-grained error approaches can be used for increased accessibility and semi-supervised labeling, as the methods model plausible explanations (some of them by testing the alignment of scores aggregated from explanations with human MQM scores). The usage for metric diagnosis is limited, as only samples are explained, and explanations have not been tested for faithfulness. The current approaches cannot be used to detect social biases.
Audiences: Flagged translation errors are simple to understand, such that these techniques could be used by non-expert and expert users. Metric developers might use them to create new data sets to train new metrics.
Limitations: The current approaches (besides Xu et al. 2023) do not have the goal to provide explanations that make the sentence-level scores more descriptive. Instead, they only search to build sentence-level metrics from LLM
outputs. As such, there might be open opportunities in leveraging the explanations provided by these models to explain to a human explainee. Another limitation are missing evaluations of faithfulness.
@@
Perturbation robustness (Row 4 in Figure 4) The methods described up to now addressed decision-understanding techniques Gehrmann et al. (2020). This paragraph and the next describe model-understanding techniques (we describe these terms in §2.2). Many recent works evaluate the robustness of MT metrics (Sai et al. 2021; Karpinska et al. 2022; He et al. 2022; Sun et al. 2022; Chen and Eger 2023; Vu et al. 2022). Robustness evaluation qualifies as model-understanding, as it usually infers properties that describe a model (see §2.2). All works in this category are structured into “black box” in Figure 4.
@@
Sai et al. (2021) propose perturbation checklists, following Ribeiro et al. (2020), that evaluate how susceptible metrics are towards predefined types of perturbations as evaluation criterion. This allows developers to check whether all invariances that are required for a specific task are fulfilled. In particular, they compare the change in the metric’s score with the change in score a human would assign after the perturbation. These templates encompass dropping or adding of context as well as negations. They show that perturbation checklists allow to pick metrics that are strong with respect to specific properties. For example, their experiments indicate MoverScore (Zhao et al. 2019) would capture fluency better than BERTScore (Zhang et al. 2020a) due to its ability to cope with the jumbling of words. The checklists allow for more fine-grained assessment of automatic evaluation metrics, which exposes their limitations and supports the design of better metrics. Chen and Eger (2023) create a similar, but fully automatic variant of this evaluation. They design several perturbation types (framed as adversarial attacks) such as addition, omission and negation. Then they measure whether a tested metric prefers a meaning preserving, but lexically dissimilar paraphrase of a source/reference sentence over a minimally attacked but lexically similar version of the source/reference sentence. This means they check whether the metric score of the attacked example is successfully lower than the metric score of a paraphrase. One of their findings is that standard metrics often have problems in detecting wrong names or numbers in translations. Karpinska et al. (2022) create DEMETR, a data set of machine translations that were perturbed using several perturbation types. They propose to test a metric’s preference between these perturbations and the original translation. Their task is easier for a metric to compute than Chen and Eger (2023), as the preference is checked directly between attacked and non-attacked version, while Chen and Eger (2023) use a dissimilar paraphrase instead of the non-attacked version. Also, they define more perturbation types than Chen and Eger (2023), but require manual annotation for some of them. He et al. (2022) also compare the original translation with certain perturbation types. They differ from the other two papers in terms of domains they are applying their method to and in the perturbation types that they propose. Vu et al. (2022) use character-level adversarial attacks to evaluate the robustness of several BERTScore variants.
@@
The WMT21 (Freitag et al. 2021b) and WMT22 (Freitag et al. 2022) metrics shared tasks introduce a similar setting for robustness tests. In WMT21, the organizers provided the participants with a challenge set composed of translations that were perturbed in a specific way, to test whether metrics reduce their score as expected when presented with respective phenomena. In WMT22, participants of a challenge set subtask (Alves et al. 2022; Avramidis and Macketanz 2022; Chen et al. 2022; Amrhein et al. 2022) designed data sets posing challenges in a similar manner. In the example presented in Figure 5, we show tests similar to those performed by techniques of this category. The metric score should be decreased if numbers or word order used in the hypothesis change.
@@
Goals: The presented perturbation robustness approaches give detailed insights into metric weaknesses. Therefore, they are most suited to be used for metric diagnosis and improvement.
Audiences: Metric developers can use these methods for diagnosis. They can also be used by metric users to check if a metric fits their use case. Regulatory entities might test if metrics fulfill their guidelines, e.g. exhibit no social biases. Lastly, affected users might gather evidence when treated unfairly.
Limitations: A drawback of the perturbation-based approaches is that they require predefined error types. Therefore, errors lacking a definition in the evaluation scheme that would fool the machine translation metrics could be missed. These approaches are faithful to some degree, namely in the domains they have been evaluated in. For small perturbation sets it could be questionable if the model generally behaves as predicted. Therefore we would suggest evaluating each property on a large, general data sample. Still, it could be advisable to employ faithfulness tests, if possible, to achieve guarantees.
@@
Linguistic properties (Row 5 in Figure 4) Here, we summarize model-understanding techniques that directly summarize various linguistic properties a metric fulfills. So far, the only method that falls into this category for MT is Kaster et al. (2021) (see linguistic properties/black box in Figure 4). They propose a model-understanding technique that decomposes the score of sentence-level BERT-based metrics into linguistic factors. In particular, they explore the degree to which metrics consider the properties syntax, semantics, morphology and lexical overlap. To do so, they measure these properties in a given data set and learn regressors that explain how much each distinct property contributes to the original metric’s score. In their experiments, they show that each metric captures semantic similarity and lexical overlap to some degree. Syntactic and morphological similarity are either captured to a smaller extent or are even negatively correlated with the metric score. Especially, MoverScore (Zhao et al. 2019) and BERTScore (Zhang et al. 2020a) have a comparatively high coefficient for the lexical score.
@@
As shown in Figure 5, these techniques might give explanations like “the metric puts a high focus on lexical properties (green bar) and a low focus on semantics (red bar)“.
@@
Note that the explanation of linguistic properties is close to the decision-understanding of quality aspects. If these are collected over many samples and a regression is performed, similar explanations might be achieved. A related approach used for another NLP task is by Opitz and Frank (2022), who retrain sentence embeddings to decompose into sub-components that match the aspects of abstract meaning representation, a format to represent the meaning of sentences. This allows them to attribute parts of the score to different aspects of a sentence. They evaluate their approach on the model-understanding level. When this would be applied to MT, it would also count towards this explanation type.
@@
We also list Kaster et al. (2021); Opitz and Frank (2022) with quality aspect explanations, as the metrics that measure the linguistic factors can also be computed on a sentence level.
Goals: The discussed linguistic properties approaches evaluate general metric properties and can be used for metric debugging. As they generate sentence/metric-level property scores, the techniques
are
not usable for word-level semi-automatic labeling. When used on a sentence-level, the techniques can aid the goal of making metrics more accessible. While not employed here, the techniques might be usable to detect social biases by introducing a set of bias detection scores from which to learn the regressors.
Audiences: The linguistic properties can be easily explained to non-experts, however metric selection should not solely be performed based on linguistic properties
and
is advised to be handled by domain experts. The methods are especially useful to debug metrics, i.e. for metric developers.
Limitations: The approach by Kaster et al. (2021) could not explain reference-free metrics well, so plausibly requires alternate explanatory variables.
The search for regressors may be inspired by quality aspect approaches (described earlier this section) where not a global metric score is reported but several scores (such as coherence, fluency, etc.). These could then compose a global MT score.
Further, some of the “property metrics” could be considered black box variables themselves and future work might replace them by more transparent factors.
One might also explore the collinearity of the different regressors.
For Kaster et al. (2021), faithfulness is likely given to some degree, around the samples that was trained on. Still, it could be advisable to employ faithfulness tests, if possible, to achieve guarantees.

### v2

@@ 4 Taxonomy
Two shared tasks for the explainability of reference-free metrics have originated from the setting in Fomicheva et al. (2022) so far (Fomicheva et al., 2021; Zerva et al., 2022). As the respective papers already give an overview of the participating methods, we will only briefly describe the methods and set our focus to describe the model access they require. We describe the model access levels of Figure 3 from left to right. Approaches that work directly on black box models are explored in Fomicheva et al. (2022), Eksi et al. (2021) and Treviso et al. (2021). They leverage model agnostic explainers such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017). These methods check how a metric’s sentence-level score changes when the input sentence is perturbed at selected positions. While the results correlate with the human annotations to some degree, techniques that use model access performed better in the shared tasks. Gradient based methods are explored by Fomicheva et al. (2022), Eksi et al. (2021), Treviso et al. (2021) and Rei et al. (2022a). These papers use methods like Integrated Gradients (Sundararajan et al., 2017) to determine the effect each input token has on the output. Attention values perform especially good for Treviso et al. (2021) and Rei et al. (2022b), leading to winning techniques for most language pairs in both shared tasks, i.e., producing the most plausible explanations. In specific, Treviso et al. (2021) succeed by scaling attention weights by the norm of value vectors and filtering for attention heads that have the highest scores in detecting errors. Rei et al. (2022b) build on their approach and (1) instead scale attention weights with the l2-norm of the gradient of value vectors and (2) designing an ensemble method for attention heads. In their final submission, they ensemble the best performing attention heads over all layers.
Word-level contextualized embeddings are used by Leiter (2021), Tao et al. (2022) and Azadi et al. (2022). Specifically, they leverage the word-level embedding distances that are inherently computed in token-matching metrics (see §2.1) and use them as importance scores. For model-specific works, Rubino et al. (2021) jointly fine-tune their metrics on the word-level and sentence-level using synthetic data, to directly output the computed word-level scores.121212Rubino et al. (2021) won the constrained track of Eval4NLP21, which did not allow for word-level finetuning. Kabir and Carpuat (2021) use LIME with Transquest (Ranasinghe et al., 2020b) and merge the word-level importance scores with additional word-level scores obtained from a pure word-level model.
After the shared tasks, Fernandes et al. (2022b) enhance explainability techniques (in case of MT they use a gradient technique) by adding learnable parameters and employing a student teacher mechanism to improve the correlation to human word-level scores.
Rei et al. (2023b) extend the work of Rei et al. (2022b) by using reference-based metrics and importance scores of the references. They construct a data set from the WMT21 MQM annotations Freitag et al. (2021b) by assuming that feature attribution scores link to annotated error spans. Evaluations on this data set show that the winning explainability techniques of the shared task perform best and that references can improve the plausibility. Notably, they also show that explanations can help to detect critical errors in translations. Based on this, they employ the explanations for metric diagnosis, to test whether their metrics are more prone to ignoring certain types of critical errors. In their limitations, they note that the explainability techniques they use have shown to be faithful in some scenarios, but also note that they did not specifically test whether they are faithful for their architecture.
Finally, we note that many more metrics can provide feature-importance like explanations by design. Similar to the base metrics used by Leiter (2021), Tao et al. (2022) and Azadi et al. (2022), most unsupervised matching metrics could be used to determine importance scores based on word- or embedding-similarities. Further, the per-word generation probabilities in text generation metrics like PRISM or BARTScore can also be interpreted as feature importance (they can also be interpreted as the confidence of the generation models; Fomicheva et al. 2020).
@@
In the example presented in Figure 4, high feature importance scores are assigned to words that were translated correctly, i.e., “Things/complicated” and lower scores indicate words that are wrong.
@@
Fine-grained errors (Row 3 in Figure 3) Some recent works detect errors on a word-level and leverage their severity, number and type to compute a sentence-level score (Lu et al., 2023a, b; Bao et al., 2023; Xu et al., 2023; Fernandes et al., 2023; Kocmi and Federmann, 2023a). This follows the ideas of the MQM annotation scheme (Lommel et al., 2014; Freitag et al., 2021a) in which human translation experts evaluate translations in a similar manner. Freitag et al. (2021a) have shown that this annotation scheme improves over previous annotation methods such as direct assessments (Graham et al., 2015). While the main goal of these works is to provide sentence-level scores, they are also explainable-by-design and the generated error-labels carry explanatory value (model-specific in Figure 3). In specific, BARTScore++ (Lu et al., 2023a) distinguishes explicit errors that are easy to detect and implicit errors. They detect the explicit errors based on BART generation probabilities. Then they generate a sentence-level score based on an additional improved hypothesis sentence in multiple BARTScore calculations.
Lu et al. (2023b) also combine fine-grained errors into a score. They first request ChatGPT to generate explanations based on MQM annotations (erroneous words, error type and error majority). In a second step, they ask a model to calculate a score from these. This is a form of chain-of-thought prompting (Wei et al., 2022), a prompting technique where a language model is instructed to first generate an explanation/reasoning before returning a result. Lu et al. (2023b) find that their approach is competitive to other state-of-the-art metrics, especially with newer versions of ChatGPT.
Xu et al. (2023) fine-tune InstructScore, a LLaMA model (Touvron et al., 2023) that generates a text containing error majority, error spans, error types (MQM) and an explanation for every error in an input sentence. The generated texts are similar to Lu et al. (2023b) but add more details on each error. Similarly, Xu et al. (2023) construct a score from the count and severity of errors. They obtain synthetic training data from GPT4 (OpenAI, 2023b), analyze failure modes of InstructScore during training and later optimize the metric on data sets with real machine translation outputs and annotated ground truth scores. Their model shows state-of-the-art results for reference-free sentence-level evaluations. Further, the detailed error annotations can be used as fine-grained explanations. Contrary to Instructscore, instead of fine-tuning, Fernandes et al. (2023) propose AutoMQM a prompting technique that leverages few-shot learning (Brown et al., 2020) and chain-of-thought prompting to generate MQM error annotations. They use it to prompt PALM 2 (Anil et al., 2023) and achieve competitive results for system- and sentence-level evaluation. In a similar way, Kocmi and Federmann (2023a) design the GEMBA-MQM prompting technique and prompt GPT4 in a few-shot and chain-of-thought setting to produce MQM annotations. A main distinction is that AutoMQM dynamically selects few-shot examples from pools of example translations and ratings based on the input. In contrast, GEMBA-MQM uses only a few, pre-selected examples and is therefore easier to apply to further language pairs. Both, AutoMQM and GEMBA-MQM calculate their final scores based on the detected errors with MQM heuristics (Freitag et al., 2021a).
The WMT 2023 shared task on quality estimation (Blain et al., 2023) also includes a new sub-track where the submitted metrics should detect fine-grained errors and their severity on MQM data, to facilitate the development of more explainable metrics. The participants explore two main directions: (1) detection with GPT4 (Rei et al., 2023a)141414Another participating team KUNMT did not submit a paper. and (2) fine-tuning models for error span detection, either based on word-level supervision (Geng et al., 2023; Li et al., 2023b) or with multi-task objectives (Rei et al., 2023a). Notably the two winning submissions from Geng et al. (2023) and Rei et al. (2023a) are ensemble approaches, where Rei et al. (2023a) also use pseudo-references from Deepl and Google Translate (if their quality is better than a threshold, as measured by another metric) (Blain et al., 2023).
Similarly, but not related to the shared task, Bao et al. (2023) fine-tune a seq2seq model to predict addition and omission errors on a word-level.
In the example presented in Figure 4, a tag “sense” flags the wrong word sense of “became”. Notably, all fine-grained error approaches we describe are very recent, which is potentially caused by the adaptation of MQM in the WMT metrics shared tasks and the recent successes of LLMs.
@@
Perturbation robustness (Row 4 in Figure 3) The methods described up to now addressed decision-understanding techniques Gehrmann et al. (2020). This paragraph and the next describe model-understanding techniques (we describe these terms in §2.2). Many recent works evaluate the robustness of MT metrics (Sai et al., 2021; Karpinska et al., 2022; He et al., 2023; Sun et al., 2022; Chen and Eger, 2023; Vu et al., 2022). Robustness evaluation qualifies as model-understanding, as it usually infers properties that describe a model (see §2.2). All works in this category are structured into “black box” in Figure 3.
Sai et al. (2021) propose perturbation checklists, following Ribeiro et al. (2020), that evaluate how susceptible metrics are towards predefined types of perturbations as evaluation criterion. This allows developers to check whether all invariances that are required for a specific task are fulfilled.
In particular, they compare the change in the metric’s score with the change in score a human would assign after the perturbation.
These templates encompass dropping or adding of context as well as negations. They show that perturbation checklists allow to pick metrics that are strong with respect to specific properties. For example, their experiments indicate MoverScore (Zhao et al., 2019) would capture fluency better than BERTScore (Zhang et al., 2020a) due to its ability to cope with the jumbling of words. The checklists allow for more fine-grained assessment of automatic evaluation metrics, which exposes their limitations and supports the design of better metrics.
@@
Chen and Eger (2023) create a similar, but fully automatic variant of this evaluation. They design several perturbation types (framed as adversarial attacks) such as addition, omission and negation. Then they measure whether a tested metric prefers a meaning preserving, but lexically dissimilar paraphrase of a source/reference sentence over a minimally attacked but lexically similar version of the source/reference sentence. This means they check whether the metric score of the attacked example is successfully lower than the metric score of a paraphrase. One of their findings is that standard metrics often have problems in detecting wrong names or numbers in translations. Karpinska et al. (2022) create DEMETR, a data set of machine translations that were perturbed using several perturbation types. They propose to test a metric’s preference between these perturbations and the original translation. Their task is easier for a metric to compute than Chen and Eger (2023), as the preference is checked directly between attacked and non-attacked version, while Chen and Eger (2023) use a dissimilar paraphrase instead of the non-attacked version.
Also, they define more perturbation types than Chen and Eger (2023), but require manual annotation for some of them. He et al. (2023) also compare the original translation with certain perturbation types. They differ from the other two papers in terms of domains they are applying their method to and in the perturbation types that they propose. Vu et al. (2022) use character-level adversarial attacks to evaluate the robustness of several BERTScore variants.
@@
The WMT21 (Freitag et al., 2021b), WMT22 (Freitag et al., 2022) and WMT23 (Freitag et al., 2023) metrics shared tasks introduce a similar setting for robustness tests. In WMT21, the organizers provided the participants with a challenge set composed of translations that were perturbed in a specific way, to test whether metrics reduce their score as expected when presented with respective phenomena. In WMT22 and WMT23, participants of a challenge set subtask (Alves et al., 2022; Avramidis and Macketanz, 2022; Chen et al., 2022; Amrhein et al., 2022; Lo et al., 2023; Amrhein et al., 2023; Avramidis et al., 2023) designed data sets posing challenges in a similar manner. In the example presented in Figure 4, we show tests like those performed by techniques of this category. Here, a model-understanding explanation could be that the metric generally assigns lower scores when input sentences are subjected to numeric or word order perturbations.
@@
Linguistic properties (Row 5 in Figure 3)
@@
Here, we summarize model-understanding techniques that directly summarize various linguistic properties a metric fulfills. So far, the only method that falls into this category for MT is Kaster et al. (2021) (see linguistic properties/black box in Figure 3). They propose a model-understanding technique that decomposes the score of sentence-level BERT-based metrics into linguistic factors. In particular, they explore the degree to which metrics consider the properties syntax, semantics, morphology and lexical overlap. To do so, they measure these properties in a given data set and learn regressors that explain how much each distinct property contributes to the original metric’s score. In their experiments, they show that each metric captures semantic similarity and lexical overlap to some degree. Syntactic and morphological similarity are either captured to a smaller extent or are even negatively correlated with the metric score. Especially, MoverScore (Zhao et al., 2019) and BERTScore (Zhang et al., 2020a) have a comparatively high coefficient for the lexical score.
@@
The example in Figure 4 shows that these techniques can give explanations that highlight which properties a metric generally focuses on. In the specific example, the metric puts a high focus on lexical properties (green bar) and a low focus on semantics (red bar).
Note that the explanation of linguistic properties is close to the decision-understanding of quality aspects. If these are collected over many samples and a regression is performed, similar explanations might be achieved.
@@
A related approach used for another NLP task is by Opitz and Frank (2022), who retrain sentence embeddings to decompose into sub-components that match the aspects of abstract meaning representation, a format to represent the meaning of sentences. This allows them to attribute parts of the score to different aspects of a sentence. They evaluate their approach on the model-understanding level. When this would be applied to MT, it would also count towards this explanation type.
@@
We also list Kaster et al. (2021) and Opitz and Frank (2022) with quality aspect explanations, as the metrics that measure the linguistic factors can also be computed on a sentence-level.